OpenStudy (anonymous):

The value of an uncirculated mint-state-65 1950 Jeffererson nickel minted in Denver is 8/5 the value a 1945 nickel minted in Philadeiphia in similar condition. Together the total value of the two coins is $26. What is the value of each coin?

OpenStudy (anonymous):

1) Consider x (l950) and y (l945)the two values. a) x = 8y/5 b) x + y = 104 Now solve the system. Put x= 8y/5 in the second equation and get: 8y/5 + y = 104 ==> 8y + 5y = 5020 ==> 13y = 520 ==> y = 520/13 ==> y = 40 and x+40 = 104 ==> x= 64 OK! 2) 9x-7 <8x + 11 First add 7 to both side and get 9x < 8x + 18 Now subtract 8x to both side and get x< 18 OK!

OpenStudy (anonymous):

where are you getting the 104 from?